1-Find the equation of the line perpendicular to the line whose equation is 2y-4x=7 and which passes through the point (1,2)
2-Given 3y-4x=5 and 4y+6=3x.Are these lines parallel,perpendicular,or neither?

Answer by stanbon(75887)

1-Find the equation of the line perpendicular to the line whose equation is
2y-4x=7 and which passes through the point (1,2)
-----
Slope of the given equation::
y = 2x + (7/2)
slope = 2
-------
Slope of the perpendicular line::
slope = -1/2
------
Answer form:: y = mx + b
Solve for "b":
2 = (-1/2)1 + b
b = 2 1/2 = (5/2)
------
Answer:: y = (-1/2)x + (5/2)
=====================================
2-Given 3y-4x=5 and 4y+6=3x.Are these lines parallel,perpendicular,or neither?
3y-4x = 5 DATA:
y = (4/3)x + (5/3)
slope = 4/3 ; intercept = (5/3)
---------------------------------------------
4y+6 = 3x DATA:
y = (3/4)x - (3/2)
slope = 3/4 ; intercept = (-3/2)
------
Answer:: not perpendicular ; not parallel ; neither
-------------
Cheers,
Stan H.
------------